OpenStudy (cookiemonster18):

(5 questions please help correct my incorrect answers) 1.Simplify x/7x+x^2 A.1/7+x;where x≠-7 B.1/7x;where x≠0 C.1/7+x;where x≠0,-7 D.1/7**** 2.Simplify -12x^4/x^4+8x^5 A.-12/1+8x;where x≠-1/8**** B.-12/1+8x;where x≠-1/8,0 C.-12/9x;where x≠0 D.-12/9 3.Simplify x^2-3x-18/x+3 A.x-3 B.x-6;where x≠-3 C.x-6;where x≠6 D.1/x+3;where x≠-3**** 4.Simplify 7/2a•5/a^2 A.35/2a^2;where a≠0 B.35/2a^2**** C.12/2a^3;where a≠0 D.35/2a^3;where a≠0 5.Multiply x-5/4x+8•(12x^2+32x+16) A.(3x+2)/4(x-5) B.(x-5)(3x+2)/4**** C.(x-5)(3x+2) D.(x-5)(12x+8)

OpenStudy (cookiemonster18):

I think the answers for 1-5 are: 1.D 2.A 3.D 4.B 5.B But I'm not sure, please help

OpenStudy (rootbeer003):

You still need help?

OpenStudy (cookiemonster18):

Yes @Rootbeer003 I do very much :'( do you think you could help?

OpenStudy (rootbeer003):

Let me take a look

OpenStudy (cookiemonster18):

Okay :) thank you @Rootbeer003 I've needed help with these problems for a while now and no one came to help me yet except you.

satellite73 (satellite73):

\[\frac{ x}{7x+x^2}=\frac{x}{x(7+x)}\]

satellite73 (satellite73):

then cancel

OpenStudy (rootbeer003):

sorry i fell asleep

OpenStudy (rootbeer003):

i should take credit for helping cause i was here first

OpenStudy (cookiemonster18):

@Rootbeer003 @Satellite73 can you tell me if 1-5 are wrong and what's the correct answer :( I hate a lot of question and I spent most on the time trying to see if I could get help and rechecking if I'm correct or not. I can tell you my other theory for the questions if you like